OpenStudy (secret66):

Which square parking lot is the largest they can cover with this much asphalt?

OpenStudy (kira_yamato):

Calculate the area for all 4 choices first.

OpenStudy (secret66):

21*21=441 22*22=484 23*23-529 24*24=576 now what next?

OpenStudy (anonymous):

As you can see 24*24 covers most of the area

OpenStudy (kira_yamato):

So we have 600 sq ft of asphalt, which is enough to cover 576 (in this case the largest). So the answer is D. The idea stems from the fact that you want to have the largest area A such that A ≤ 600

OpenStudy (secret66):

Thanks Kira

OpenStudy (kira_yamato):

No problems ^_^